you are given the equation A= -3w^2+102w find when area is equal to 360
Answer by Alan3354(69443)

you are given the equation A= -3w^2+102w find when area is equal to 360
---------------
A= -3w^2+102w
360 = -3w^2 + 102w
3w^2 - 102w + 360 = 0
w^2 - 34w + 120 = 0
(w-4)*(w-30) = 0
w = 4, w = 30